OpenStudy (jjuden):

Kevin’s age can be represented by the expression 12p^3. Jake’s age can be represented by the expression 4p^5. What is the ratio of Kevin’s age to Jake’s age?

OpenStudy (mathmale):

\[\frac{ Kevin's.age }{ Jake's.age }= \frac{ Write.K's.age.here }{ Write.j's.age.here }\]

OpenStudy (mathmale):

Then simplify the right side of this equation. Hint: use a rule of exponents to do this.

OpenStudy (anonymous):

Ratio of Kevin’s age to Jake’s age = Kevin's age / Jake's age \[\huge = \frac{12p^3}{4p^4} =\frac{ 2 \times 2 \times 3 \times p \times p \times p}{2 \times 2 \times p \times p \times p \times p } \] \[\huge = \frac{12p^3}{4p^4} =\frac{ 3}{ p } = 3 : p\]

OpenStudy (mathmale):

Actually, that \[\huge Age.ratio= \frac{12p^3}{4p^4} =\frac{ 3}{ p } = 3 : p\]should be\[\huge = \frac{12p^3}{4p^5} =?\] @jjuden : Please finish this.